Homework Question 2: Find the quantities that solve the equations Q-2-.502 Answer to Homework Question 2:
Answer #1

Q​​​​​​2=Q

Q​​​​​​1=q

Equation 1

Q = 2-0.5q

Q+0.5q-2=0 ------------eq1

Equation 2

q= 2-0.5Q

0.5Q +q -2 =0 -------eq2

2*(eq 1) - (eq 2) = 0

2(Q+0.5q-2)-(0.5Q+q-2)=0

2Q+q-4-0.5Q-q+2 = 0

1.5Q-2 = 0

Q =2/1.5

Q = 1.33

q = 2-0.5(1.33) = 1.33

Answer: 1.33

For the above equation for the condition

Q1 = Q2

All the natural number from 0,1,2,3…..............N will solve the equation